# A Risk Taxonomy for Evaluating AI-Powered Psychotherapy Agents

### v1

@@ 4. Overview of the Risk Taxonomy
The refined risk taxonomy, derived from literature (Linden, 2013; Rozental et al., 2018; Klatte et al., 2023), expert interviews, and DSM-5 criteria (Black and Grant, 2014), provides a structured framework for identifying potential negative outcomes in AI psychotherapy. The full taxonomy can be found in Appendix A. It has two main categories:
1. Immediate Risk: This category encompasses situations of imminent danger to the user or others, requiring urgent intervention. Examples include explicit suicidal intent with plan and means, threats of harm to others, or severe psychological decompensation such as acute psychosis. Identification relies primarily on dialogue-based triggers. Evaluation assesses the AI’s ability to recognize these triggers and respond according to predefined, ethically-informed action plans, such as attempting de-escalation, recommending emergency resources (e.g., 911, 988 Lifeline), and contacting/flagging a human supervisor immediately.
2. Potential Risk: This category refers to emerging states or interaction patterns suggesting heightened vulnerability to negative therapeutic outcomes, functional deterioration, or adverse impacts on the therapeutic process, often aligning with unwanted events in the NEQ or UE-ATR (Rozental et al., 2019; Linden, 2013). Key subcategories are: 2.1. Symptom Exacerbation, Emergence of New Symptoms, or Maladaptive Coping, which involves worsening of clinical symptoms (e.g., increased negative thoughts or feelings) or new maladaptive behaviors (e.g., Worry/Rumination Frequency); 2.2. Patient Destabilization or Feeling Overwhelmed, indicating the user is struggling with the therapeutic process (e.g., increased confusion, emotional lability); 2.3. Therapeutic Alliance Damage or Negative Relational Dynamics, covering negative shifts in the user-AI relationship including mistrust, feeling judged, or over-dependence; and 2.4. Reduced Engagement or Potential for Premature Termination, signaling decreased participation or commitment within the therapeutic process or current therpay session.
Identification of these potential risks involves tracking dynamic changes in internal cognitive, emotional, and behavioral factors, which are conceptualized as scales (e.g., from 1-low to 10-high for intensity or frequency, such as ”Hopelessness Level” or ”Collaboration/Engagement Level”). The up (↑) and down (↓) arrows accompanying these factors in the full taxonomy (see Appendix A) specify the direction of change from a user’s baseline state that may signify an increased potential risk. For example, an increase (↑) in ”Hopelessness Level” or a decrease (↓) in ”Collaboration/Engagement Level” would be flagged as potentially problematic.
Evaluation should assess harmful shifts in these factors from baseline values, allowing for the tracking of subtle deteriorations or emerging vulnerabilities. While empirically-grounded numeric thresholds for what constitutes a significant change is a goal for future work, the current approach suggests that any major deviation beyond simple, minor fluctuations around a baseline cognitive factor value (e.g., more than +/- 1 point on a 1-10 scale from the baseline), or other abnormal fluctuations, could indicate an increased potential risk requiring attention. Furthermore, as expert participants highlighted, these indicators should ideally be monitored over the course of a session or across multiple interactions to understand their trajectory and contextual significance, rather than relying solely on isolated measurements. For instance, a consistent decrease (↓) in the Motivation/Self-Efficacy Level factor over multiple sessions could strongly suggest an increased risk for Reduced Engagement / Potential for Premature Termination.
5. Expected Use Cases and Adaptability
The taxonomy standardizes safety evaluation criteria for AI-powered psychotherapy agents. There are several ways the taxonomy can be used to improve safety. We outline several use cases here.
5.1. Monitoring Real User Interactions
@@
The risk taxonomy could be used to identify emergent risks in real-time, as a user interacts with an AI therapist, for example to provide early termination of a harmful session, or feedback to developers or regulatory bodies. Monitoring should be performed with ethical consent, and could be accomplished by tracking user risk factors from pre- to post-interaction to identify ”Potential Risk” shifts, while dialogue tools scan for ”Immediate Risk” triggers.
5.2. Evaluation with Simulated Patients
The risk taxonomy could also be used to ”monitor” interactions between an AI therapist and a simulated patient, such as (Steenstra et al., 2025), providing a powerful evaluation testbed that can be used to assess systems for safety and potential harm. During these simulated interactions, simulated patient agents’ dynamic internal states corresponding to ”Potential Risk” factors (e.g., ”Hopelessness Level) are updated. This allows systematic testing by comparing baseline to post-interaction values of internal patient factors and observing AI responses to diverse personas, a methodology previously validated in our development of SimPatient (Steenstra et al., 2025), an LLM-powered system where embodied simulated patients featured dynamic internal cognitive states (e.g., self-efficacy) that evolved during interactions.
5.3. Benchmarking and Comparative Analysis
The risk taxonomy provides a standardized framework essential for robust benchmarking and comparative analysis of different AI psychotherapist models or system versions. By systematically evaluating AI therapists against the defined risk categories and factors, consistent and comparable risk profiles can be generated for each system. This process can highlight the relative strengths and weaknesses of various AI approaches in terms of their propensity to elicit or mitigate specific types of risks. For example, one model might excel at de-escalating immediate suicidal ideation but perform poorly by increasing therapeutic alliance damage, while another might show the opposite pattern. Such comparative insights are invaluable for developers seeking to enhance model safety, for organizations making informed deployment decisions based on specific use-case risk tolerances, and for fostering a more transparent and accountable development ecosystem focused on minimizing harm.
5.4. Identifying Unexpected Outcomes
@@
Finally, the risk taxonomy serves as a crucial tool for identifying unexpected or emergent adverse outcomes that may arise during automated therapy. This is achieved through systematic monitoring of user interactions and therapeutic outcomes, continually assessed against the taxonomy’s defined risk factors and categories. Such monitoring can reveal significant deviations from established baseline values—which ideally exhibit only safe fluctuations—or flag the emergence of previously unrecognized harmful patterns. When substantial deviations or novel harmful patterns are detected, a deeper analysis is warranted. This involves scrutinizing the conversational context surrounding these occurrences to understand the specific interaction dynamics or AI behaviors that may have precipitated them. Uncovering these root causes is invaluable, as it directly informs targeted safety improvements and iterative refinements of the AI model, ultimately leading to a more robust and safer therapeutic tool.
While our risk taxonomy provides a general framework, it can also be augmented for specific populations or problems. For instance, evaluating an AI psychotherapist for helping patients with substance use disorders could add ”Potential Risk” factors such as ”Substance Craving Intensity”, ”Ambivalence about changing substance use” ”Substance Use Frequency/Amount,” ”Level of Intoxication/Withdrawal Symptoms”, and relational dynamics like ”Dishonesty/Minimization regarding substance use.” This tailoring allows for more nuanced risk assessment. Future work includes validating the taxonomy against human expert judgments.
@@ 6. Conclusion
AI psychotherapists present significant opportunities and risks. Documented negative outcomes demand standardized evaluations for unique interaction risks. We developed the risk taxonomy, from an iterative process involving literature, expert insights, and DSM-5, and harm checklists. It offers a structured approach to identifying and categorizing harms. Its overview, grounding, and adaptable use cases (tracking risk factor changes from baseline to post-interaction) aim to equip stakeholders to ensure AI psychotherapy safety, fostering innovation that prioritizes user well-being and minimizes unintended negative consequences.
Appendix A Risk Taxonomy

### v2

@@ 4. Overview of the Risk Ontology
The refined risk ontology (see Appendix A), derived from literature (e.g., (Linden, 2013; Rozental et al., 2018, 2016; Klatte et al., 2023)), expert interviews, and DSM-5 criteria (Association et al., 2013), offers a structured framework to identify potential negative outcomes in AI psychotherapy. It has three overarching categories:
1. Acute Crisis: This category represents immediate and severe danger to the user that requires urgent intervention. Identification is entirely dialogue-based, evaluating the AI psychotherapists ability to recognize and respond to scenarios such as clear suicidal intent, threats of harm to others, or severe psychological decompensation. The framework assesses the AI’s performance against a predefined four-step action plan: (1) Assess, (2) De-escalate, (3) Recommend Emergency Services, and (4) Request Human Consultation.
2. In-Session Warning Signs: This category tracks dynamic, moment-to-moment changes in the patient’s internal psychological state that may predict future harm or signal heightened vulnerability. These are not direct harms but are leading indicators. The framework operationalizes key psychological constructs (e.g., Hopelessness, Self-Efficacy) as internal state variables, which are conceptualized as a 5-point Likert scale. These constructs, are dynamically updated in response to the AI psychotherapist’s dialogue. Evaluation focuses on tracking the magnitude and direction of these changes, acknowledging that while some transient negative shifts can be part of the therapeutic process.
3. Potential Real-World Consequences: This category represents the tangible, adverse outcomes that could manifest in the patient’s life following a therapeutic interaction, often aligning with NEQ and UE-ATR unwanted events (Rozental et al., 2016; Linden, 2013). Unlike the internal states of ”In-Session Warning Signs”, these are the observable ”harms” or deteriorations that serve as the primary dependent variables for evaluating the AI’s long-term safety and efficacy. Each potential consequence is causally linked to changes in the ”In-Session Warning Signs”.
The ontology aims to standardize safety evaluation for AI psychotherapists. We outline several use cases for improving safety:
@@
Monitoring Real User Interactions: The ontology can identify emergent risks in real-time during user-AI psychotherapist interactions, enabling early termination of harmful sessions or providing feedback to developers and regulators. Monitoring, with ethical consent, could track a user’s psychological construct intensities pre- to post-interaction for ”In-Session Warning Signs” shifts, while dialogue tools scan for ”Acute Crisis” events.
Evaluation with Simulated Patients: The risk ontology can also structure the monitoring of interactions between an AI psychotherapist and a simulated patient. This approach is exemplified in our work on SimPatient, a testbed designed for safety and quality of care assessment (Steenstra et al., 2025). Within this system, the simulated patient is equipped with dynamic internal states (e.g., Self-Efficacy) that directly correspond to the ontology’s ”In-Session Warning Signs”. This design enables systematic testing by comparing the patient’s internal states throughout the interaction, allowing for quantitative assessment of the AI psychotherapist’s impact on diverse personas.
Benchmarking and Comparative Analysis: The ontology provides a standardized framework for robust benchmarking and comparing AI psychotherapist models and versions. Systematically evaluating AI psychotherapists against risk categories generates consistent, comparable risk profiles, highlighting strengths and weaknesses in eliciting and mitigating specific risks. Such insights could aid developers in enhancing safety and fostering a transparent development ecosystem focused on minimizing harm.
@@
Identifying Unexpected Outcomes: The ontology is crucial for identifying unexpected and emergent adverse outcomes in automated therapy. Systematic monitoring of user interactions and outcomes against ontology categories can reveal significant deviations from baselines or flag harmful patterns. Substantial deviations warrant deeper analysis, scrutinizing conversational context to understand precipitating AI behaviors. Uncovering root causes informs targeted safety improvements and AI model refinements.
Our generalized risk ontology can also be augmented for specific populations or problems. For instance, evaluating an AI psychotherapist for patients with substance use disorders, could add a ”In-Session Warning Signs” factor like ”Substance Craving Intensity”. This tailoring enables nuanced risk assessment.
A critical next step is the formal validation of the ontology itself. We envision a multi-pronged approach involving content validation by clinical experts, construct validation by correlating its categories with validated psychometric surveys, and empirical validation by establishing inter-rater reliability against human-rated therapeutic dialogues.
@@ 5. Conclusion
AI psychotherapists present significant opportunities and risks. Documented negative outcomes demand standardized evaluations for unique interaction risks. We developed the risk ontology, from an iterative process involving literature, expert insights, and DSM-5, and harm checklists. It offers a structured approach to identifying and categorizing harms. Its overview, grounding, and adaptable use cases (tracking in-session warning signs) aim to equip stakeholders to ensure AI psychotherapy safety, fostering innovation that prioritizes user well-being and minimizes unintended negative consequences.
Appendix A Supplementary Material – AI Therapy Risk Ontology
Figure 1. A High-Level Overview of the AI Therapy Risk Ontology:
Figure 2. Action Plans for Responding to Acute Crisis Scenarios:
Figure 3. Definitions of Psychological Constructs Used as In-Session Warning Signs:
Figure 4. Potential Behavioral Harms and Their Predictive Psychological Constructs:
Figure 5. Potential Cognitive and Affective Harms and Their Predictive Psychological Constructs:
Figure 6. Potential Relational Harm and its Predictive Psychological Constructs: